Part A Capital Budgeting (Show all workings 50 marks)

Background:

(EV) GOGreen Motors is considering a new project to produce electric vehicles for the Australian domestic market and international markets. GOGreen has identified a property/plant that was formerly used to build petrol fueled motor vehicles that could be refitted at minimal cost to manufacture the new EV's.   

GOGreen is targeting Australian metrolpolitan centres for initial sales and expanding into regional centres over the next five years. International demand for EVs is being driven by China and GOGreen has been in negotiation to provide vehicles to the Chinese market in 2020. Problem: GOGreen has made the following projections: • In the first year 2,000 units will be sold, growing at 10% per annum. • The price for each unit in the first year will be AU$50,000. This price will increase each year by 5%. • Variable costs are 60% of the sales price, which will grow by 3% each year. • Fixed costs are $5 mil pa, which are expected to grow by 2% each year. • The project is for a term of 5 years. The projected growth of the EV line is expected to outgrow the plant at this time, hence the plan will be sold at the end of 5 years. • Initial investment into manufacturing equipment of $100 million;equipment may be depreciated at 20% straight-line (prime cost) method. • In 5 years, the plant will be worth 10% of its' purchase price. • Working capital $3 million. • GOGreen's required rate of return is 4.5%. • The tax rate for GOGreen is 30%. (a) Prepare an excel spreadsheet calculating: • After-tax cash flows (in table format) (15 marks) • Payback period (3 marks) • Net present value (5 marks) • Profitability index (2 marks) (b) You are asked to present a report on your findings regarding the upgrade proposal. Make a recommendation to Management on whether they should proceed with the project or not. Explain the criteria on which you have based your decision. (10 marks) (c) It has come to your attention that variable costs are anticipated to rise by 10% per annum due to the prospective growth within the industry. Would you recommend to proceed with the project? (Show all calculations). (5 marks) (d) You have been asked to provide a further evaluation regarding the alternative use of the plant for the purpose of manufacturing electric buses, however the project life will be for 10 years. Explain how financial managers may evaluate both projects that are of unequal lives. (10 marks)
Answer #1

a]

Operating cash flow (OCF) each year = income after tax + depreciation

At the end of year 5, the equipment is sold. As it is fully depreciated, the entire sale proceeds will be subject to tax as the book value is zero

The working capital is assumed to be recovered at the end of year 5

NPV is calculated using NPV function in Excel

NPV is $90,886,337

Payback period is the time taken for the cumulative cash flows to equal zero. Payback period = 2 + (36,806,000 / 41,825,412) = 2.88 years

PI = (NPV + initial investment) / initial investment = 90,886,337 / (90,886,337 + 103,000,000) = 1.88

b]

Management should proceed with the project as the NPV is positive and the PI is more than 1

c]

With the variable costs increasing at 10%, the NPV is reduced to $53,407,747.

The management should still proceed with the project as the NPV is still positive

d]

Projects with unequal lives can be compared by using equivalent annual annuity (EAA) method.

EAA = (NPV * r) / (1 - (1 + r)-n), where

r = required rate of return

n = project life

The project with higher EAA should be chosen